OpenStudy (anonymous):

i need help!!! i just finished a parabola graph now on the same axes i have to graph y=2x-3. I know its supposed to be a line graph

OpenStudy (mertsj):

So mark -3 on the y axis. From that point, which is (0,-3) go up 2 units and then 1 unit to the right. Since two points determine a line, draw the line.

OpenStudy (anonymous):

so this is rise over run

OpenStudy (mertsj):

exactly. The slope is 2 or 2 over 1 which is rise over run

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

how did you get one

OpenStudy (mertsj):

The denominator of all whole numbers, when written in rational number form, is 1

OpenStudy (mertsj):

2=2/1 3=3/1 and so on.

OpenStudy (anonymous):

example: y= x-1 rise 1 up and run 1 to the left

OpenStudy (mertsj):

no. The coefficient of x in your example is 1 which is 1/1 so rise one and go one to the right.

OpenStudy (anonymous):

but isn't the one negative

OpenStudy (mertsj):

The y intercept is the -1 on the end. So the line crosses the y axis at (0,-1) From that point, follow the directions given by the slope which is 1/1 that is, rise 1 and go right 1

OpenStudy (mertsj):



OpenStudy (mertsj):